Suppose a market research company finds that at a price of p = $30, they would sell x = 45 tiles each month. If they lower the price to p = $20, then more people would purchase the tile, and they can expect to sell x = 60 tiles in a month’s time. Find the equation of the line for the demand equation. Write your answer in the form p = mx + b. Hint #1: Write an equation using two points in the form (x, p). Hint #2: The slope formula is needed.
Answer by stanbon(75887)

Suppose a market research company finds that at a price of p = $30, they would sell x = 45 tiles each month. If they lower the price to p = $20, then more people would purchase the tile, and they can expect to sell x = 60 tiles in a month’s time. Find the equation of the line for the demand equation. Write your answer in the form p = mx + b. Hint #1: Write an equation using two points in the form (x, p). Hint #2: The slope formula is needed.
price = m*quantity + b
Two points:
(45,30) and (20,60)
------
slope = (60-30)/(20-45) = 30/-25 = -6/5
------
p = mx + b
Solve for "b":
60 = (-6/5)*20 = b
b = 60 + 24
b = 84
---
Equation:
p(x) = (-6/5)x + 84
=========================
Cheers,
Stan H.
========================